# t-DGR: A Trajectory-Based Deep Generative Replay Method for Continual Learning in Decision Making

### v1

@@ Abstract
Deep generative replay has emerged as a promising approach for continual learning in decision-making tasks. This approach addresses the problem of catastrophic forgetting by leveraging the generation of trajectories from previously encountered tasks to augment the current dataset. However, existing deep generative replay methods for continual learning rely on autoregressive models, which suffer from compounding errors in the generated trajectories. In this paper, we propose a simple, scalable, and non-autoregressive method for continual learning in decision-making tasks using a generative model that generates task samples conditioned on the trajectory timestep. We evaluate our method on Continual World benchmarks and find that our approach achieves state-of-the-art performance on the average success rate metric among continual learning methods. Code is available at https://github.com/WilliamYue37/t-DGR.
@@ 1 Introduction
Continual learning, also known as lifelong learning, is a critical challenge in the advancement of general artificial intelligence, as it enables models to learn from a continuous stream of data encompassing various tasks, rather than having access to all data at once [1]. However, a major challenge in continual learning is the phenomenon of catastrophic forgetting, where previously learned skills are lost when attempting to learn new tasks [2].
@@
In the field of cognitive neuroscience, the Complementary Learning Systems theory offers insights into how the human brain manages memory. This theory suggests that the brain employs two complementary learning systems: a fast-learning episodic system and a slow-learning semantic system [3, 4, 5]. The hippocampus serves as the episodic system, responsible for storing specific memories of unique events, while the neocortex functions as the semantic system, extracting general knowledge from episodic memories and organizing it into abstract representations [6].
@@
Drawing inspiration from the human brain, deep generative replay (DGR) addresses the catastrophic forgetting issue in decision-making tasks by using a generative model as the hippocampus to generate trajectories from past tasks and replay them to the learner which acts as the neocortex (Figure 2) [7]. The time-series nature of trajectories in decision-making tasks sets it apart from continual supervised learning, as each timestep of the trajectory requires sufficient replay. In supervised learning, the learner’s performance is not significantly affected if it performs poorly on a small subset of the data. However, in decision-making tasks, poor performance on any part of the trajectory can severely impact the overall performance. Therefore, it is crucial to generate state-action pairs that accurately represent the distribution found in trajectories. Furthermore, the high-dimensional distribution space of trajectories makes it computationally infeasible to generate complete trajectories all at once.
@@
To address the issues in current DGR methods, we propose a simple, scalable, and non-autoregressive trajectory-based DGR method. We define a generated trajectory as temporally coherent if the transitions from one state to the next appear realistic (refer to Section 3.4 for a formal definition). Given that current decision-making methods are trained on state-action pairs, we do not require trajectories to exhibit temporal coherence. Instead, our focus is on ensuring an equal number of samples generated at each timestep of the trajectory to accurately represent the distribution found in trajectories. To achieve equal sample coverage at each timestep, we train our generator to produce state observations conditioned on the trajectory timestep, and then sample from the generator conditioned on each timestep of the trajectory. The intuition behind our method is illustrated in Figure 1.
To evaluate the effectiveness of our proposed method, t-DGR, we conducted experiments on the Continual World benchmarks CW10 and CW20 [8] using imitation learning. Our results indicate that t-DGR achieves state-of-the-art performance in terms of average success rate when compared to other top continual learning methods.
@@ 2 Related Work
This section provides an overview of existing continual learning methods within the context of “General Continual Learning", with a particular focus on pseudo-rehearsal methods.
@@ 2.1 Continual Learning in the Real World
As the field of continual learning continues to grow, there is an increasing emphasis on developing methods that can be effectively applied in real-world scenarios [9, 10, 11, 12, 13]. The concept of “General Continual Learning" was introduced by Buzzega et al. [14] to address certain properties of the real world that are often overlooked or ignored by existing continual learning methods. Specifically, two important properties, bounded memory and blurry task boundaries, are emphasized in this work. Bounded memory refers to the requirement that the memory footprint of a continual learning method should remain bounded throughout the entire lifespan of the learning agent. This property is crucial to ensure practicality and efficiency in real-world scenarios. Additionally, blurry task boundaries highlight the challenge of training on tasks that are intertwined, without clear delineation of when one task ends and another begins. Many existing methods fail to account for this characteristic, which is common in real-world learning scenarios. While there are other significant properties associated with continual learning in the real world, this study focuses on the often-neglected aspects of bounded memory and blurry task boundaries. By addressing these properties, we aim to develop methods that are more robust and applicable in practical settings.
@@ 2.2 Continual Learning Methods
Regularization methods in continual learning focus on incorporating constraints during model training to promote the retention of past knowledge. One simple approach is to include an $L_{2}$ penalty in the loss function. Elastic Weight Consolidation (EWC) builds upon this idea by assigning weights to parameters based on their importance for previous tasks using the Fisher information matrix [15]. MAS measures the sensitivity of parameter changes on the model’s output, prioritizing the retention of parameters with a larger effect [16]. VCL leverages variational inference to minimize the Kullback-Leibler divergence between the current and prior parameter distributions [17]. Progress and Compress learns new tasks using a separate model and subsequently distills this knowledge into the main model while safeguarding the previously acquired knowledge [18]. However, regularization methods may struggle with blurry task boundaries as they rely on knowledge of task endpoints to apply regularization techniques effectively. In our experiments, EWC was chosen as the representative regularization method based on its performance in the original Continual World experiments [8].
@@
Architecture-based methods aim to maintain distinct sets of parameters for each task, ensuring that future learning does not interfere with the knowledge acquired from previous tasks. Packnet [19], UCL [20], and AGS-CL [21] all safeguard previous task information in a neural network by identifying important parameters and freeing up less important parameters for future learning. Identification of important parameters can be done through iterative pruning (Packnet), parameter uncertainty (UCL), and activation value (AGS-CL). However, a drawback of parameter isolation methods is that each task requires its own set of parameters, which may eventually exhaust the available parameters for new tasks and necessitate a dynamically expanding network without bounded memory [22]. Additionally, parameter isolation methods require training on a single task at a time to prune and isolate parameters, preventing concurrent learning from multiple interwoven tasks. In our experiments, PackNet was selected as the representative architecture-based method based on its performance in the original Continual World experiments [8].
@@
Pseudo-rehearsal methods mitigate the forgetting of previous tasks by generating synthetic samples from past tasks and replaying them to the learner. Deep generative replay (DGR) (Figure 2) utilizes a generative model, such as generative adversarial networks [23], variational autoencoders [24], or diffusion models [25, 26], to generate the synthetic samples. Originally, deep generative replay was proposed to address continual supervised learning problems, where the generator only needed to generate single data point samples [7]. However, in decision-making tasks, expert demonstrations consist of trajectories (time-series) with a significantly higher-dimensional distribution space.
One existing DGR method generates individual state observations i.i.d. instead of entire trajectories. However, this approach leads to a higher sample complexity compared to generating entire trajectories. The sample complexity of generating enough individual state observations i.i.d. to cover every portion of the trajectory $m$ times can be described using the Double Dixie Cup problem [27]. For trajectories of length $n$, it takes an average of $\Theta(n\log n+mn\log\log n)$ i.i.d. samples to ensure at least $m$ samples for each timestep. In scenarios with limited replay time (small $m$) and long trajectories (large $n$) the sample complexity can be approximated as $\Theta(n\log n)$ using the Coupon Collector’s problem [28]. The additional $\Theta(\log n)$ factor reduces the likelihood of achieving complete sample coverage of the trajectory when the number of replays or replay time is limited, especially considering the computationally expensive nature of current generative methods. Furthermore, there is a risk that the generator assigns different probabilities to each timestep of the trajectory, leading to a selective focus on certain timesteps rather than equal representation across the trajectory.
Another existing DGR method is autoregressive trajectory generation. In the existing autoregressive method, CRIL, a generator is used to generate samples of the initial state, and a dynamics model predicts the next state based on the current state and action [29]. However, even with a dynamics model accuracy of 99% and a 1% probability of deviating from the desired trajectory, the probability of an autoregressively generated trajectory going off course is $1-0.99^{n}$, where $n$ denotes the trajectory length. With a trajectory length of $n=200$ (as used in our experiments), the probability of an autoregressively generated trajectory going off course is $0.871-0.99^{200}=$ 0.87. This example demonstrates how the issue of compounding error leads to a high probability of failure, even with a highly accurate dynamics model.
@@ 3.1 Imitation Learning
Imitation learning algorithms aim to learn a policy $\pi_{\theta}$ parameterized by $\theta$ by imitating a set of expert demonstrations $D=\{\tau_{i}\}_{i=1\ldots M}$. Each trajectory $\tau_{i}$ consists of a sequence of state-action pairs $\{(s_{j},a_{j})\}_{j=1\ldots|\tau_{i}|}$ where $|\tau_{i}|$ is the length of the trajectory. Each trajectory comes from a task $\mathcal{T}$ which is a Markov decision process that can be represented as a tuple $\langle S,A,T,\rho_{0}\rangle$ with state space $S$, action space $A$, transition dynamics $T:S\times A\times S\to[0,1]$, and initial state distribution $\rho_{0}$. Various algorithms exist for imitation learning, including behavioral cloning, GAIL [30], and inverse reinforcement learning [31]. In this work, we use behavioral cloning where the objective can be formulated as minimizing the loss function:
@@ 3.3 Diffusion Probabilistic Models
Diffusion probabilistic models [25, 26] generate data through a learned reverse denoising diffusion process $p_{\theta}(x_{t-1}\mid x_{t})$. The forward diffusion process $q(x_{t}\mid x_{t-1})$ gradually adds Gaussian noise to an input $x_{0}$ at each time step $t$, ultimately resulting in pure noise $x_{T}$ at $t=T$. The forward diffusion process is defined as:
@@
where $10<\beta_{t}<$ 1 is defined by a known variance schedule. In our implementation, we adopted the cosine schedule proposed by Nichol et al. [32]. For a sufficiently large time horizon $T$ and a well-behaved variance schedule, $x_{T}$ approximates an isotropic Gaussian distribution. If we had the reverse diffusion process $p(x_{t-1}\mid x_{t})$, we could sample $x_{T}\sim\mathcal{N}(0,\mathbf{I})$ and obtain a sample from $q(x_{0})$ by denoising $x_{T}$ with $p(x_{t-1}\mid x_{t})$. However, computing $p(x_{t-1}\mid x_{t})$ is intractable as it necessitates knowledge of the distribution of all possible $x_{t}$. Instead, we approximate $p(x_{t-1}\mid x_{t})$ using a neural network:
@@
Since $q$ and $p_{\theta}$ can be viewed as a variational auto-encoder [24], we can use the variational lower bound to minimize the negative log-likelihood of the reverse process. We can express $\mu_{\theta}(x_{t},t)$ from Equation 4 as:
@@ 4 Method
We employ a U-net [33] trained on the loss specified in Equation 6 to implement the generative diffusion model $G_{\gamma}$. Since we utilize proprioceptive observations in our experiments, $\pi_{\theta}$ is implemented with a multi-layer perceptron trained on the loss specified in Equation 1.
@@ 5.1 Experimental Setup
We evaluate our method on the Continual World benchmarks CW10 and CW20 [8], along with our own “General Continual Learning" variant of CW10 called GCL10. CW10 consists of a sequence of 10 Meta-World [34] tasks, where each task involves a Sawyer arm manipulating one or two objects in the Mujoco physics simulator. For computational efficiency, we provide the agents with proprioceptive observations. Notably, the observation and action spaces are continuous and remain consistent across all tasks. CW20 is an extension of CW10 with the tasks repeated twice. To our knowledge, Continual World is the only standard continual learning benchmark for decision-making tasks. GCL10 gives data to the learner in 10 sequential buckets $B_{1},\ldots,B_{10}$. Data from task $\mathcal{T}_{i}$ from CW10 is split evenly between buckets $B_{i-1}$, $B_{i}$, and $B_{i+1}$, except for the first and last task. Task $\mathcal{T}_{1}$ is evenly split between buckets $B_{0}$ and $B_{1}$, and task $\mathcal{T}_{10}$ is evenly split between buckets $B_{9}$ and $B_{10}$.
In order to ensure bounded memory usage, we adopt a one-hot vector approach to condition the model on the task, rather than maintaining a separate final neural network layer for each individual task. Additionally, we do not allow separate biases for each task, as originally done in EWC [15]. Expert demonstrations for training are acquired by gathering 100 trajectories per task using hand-designed policies from Meta-World, with each trajectory limited to a maximum of 200 steps. Importantly, the learner model remains consistent across different methods and benchmark evaluations. Moreover, we maintain a consistent replay ratio of $r=0.9$ across all pseudo-rehearsal methods.

### v2

@@ Abstract
Deep generative replay has emerged as a promising approach for continual learning in decision-making tasks. This approach addresses the problem of catastrophic forgetting by leveraging the generation of trajectories from previously encountered tasks to augment the current dataset. However, existing deep generative replay methods for continual learning rely on autoregressive models, which suffer from compounding errors in the generated trajectories. In this paper, we propose a simple, scalable, and non-autoregressive method for continual learning in decision-making tasks using a generative model that generates task samples conditioned on the trajectory timestep. We evaluate our method on Continual World benchmarks and find that our approach achieves state-of-the-art performance on the average success rate metric among continual learning methods.
@@ 1 Introduction
Continual learning, also known as lifelong learning, is a critical challenge in the advancement of general artificial intelligence, as it enables models to learn from a continuous stream of data encompassing various tasks, rather than having access to all data at once (Ring, 1994). However, a major challenge in continual learning is the phenomenon of catastrophic forgetting, where previously learned skills are lost when attempting to learn new tasks (McCloskey & Cohen, 1989).
@@
In the field of cognitive neuroscience, the Complementary Learning Systems theory offers insights into how the human brain manages memory. This theory suggests that the brain employs two complementary learning systems: a fast-learning episodic system and a slow-learning semantic system (McClelland et al., 1995; Kumaran et al., 2016; Mcclelland et al., 1995). The hippocampus serves as the episodic system, responsible for storing specific memories of unique events, while the neocortex functions as the semantic system, extracting general knowledge from episodic memories and organizing it into abstract representations (O’Reilly & Norman, 2002).
@@
Drawing inspiration from the human brain, deep generative replay (DGR) addresses the catastrophic forgetting issue in decision-making tasks by using a generative model as the hippocampus to generate trajectories from past tasks and replay them to the learner which acts as the neocortex (Figure 2) (Shin et al., 2017). The time-series nature of trajectories in decision-making tasks sets it apart from continual supervised learning, as each timestep of the trajectory requires sufficient replay. In supervised learning, the learner’s performance is not significantly affected if it performs poorly on a small subset of the data. However, in decision-making tasks, poor performance on any part of the trajectory can severely impact the overall performance. Therefore, it is crucial to generate state-action pairs that accurately represent the distribution found in trajectories. Furthermore, the high-dimensional distribution space of trajectories makes it computationally infeasible to generate complete trajectories all at once.
@@
To address the issues in current DGR methods, we propose a simple, scalable, and non-autoregressive trajectory-based DGR method. We define a generated trajectory as temporally coherent if the transitions from one state to the next appear realistic (refer to Section 3.4 for a formal definition). Given that imitation learning methods and many reinforcement learning methods are trained on single transition samples (e.g. state-action pairs), we do not require trajectories to exhibit temporal coherence. Instead, our focus is on ensuring an equal number of samples generated at each timestep of the trajectory to accurately represent the distribution found in trajectories. To achieve equal sample coverage at each timestep, we train our generator to produce state observations conditioned on the trajectory timestep, and then sample from the generator conditioned on each timestep of the trajectory. The intuition behind our method is illustrated in Figure 1.
To evaluate the effectiveness of our proposed method, t-DGR, we conducted experiments on the Continual World benchmarks CW10 and CW20 (Wolczyk et al., 2021) using imitation learning. Our results indicate that t-DGR achieves state-of-the-art performance in terms of average success rate when compared to other top continual learning methods.
@@ 2 Related Work
This section provides an overview of existing continual learning methods, with a particular focus on pseudo-rehearsal methods.
@@ 2.1 Continual Learning in the Real World
As the field of continual learning continues to grow, there is an increasing emphasis on developing methods that can be effectively applied in real-world scenarios (Wang et al., 2024; Aljundi et al., 2019; Bang et al., 2021; Hsu et al., 2019; van de Ven & Tolias, 2019). The concept of “General Continual Learning” was introduced by Buzzega et al. (2020) to address certain properties of the real world that are often overlooked or ignored by existing continual learning methods. Specifically, two important properties, bounded memory and blurry task boundaries, are emphasized in this work. Bounded memory refers to the requirement that the memory footprint of a continual learning method should remain bounded throughout the entire lifespan of the learning agent. This property is crucial to ensure practicality and efficiency in real-world scenarios. Additionally, blurry task boundaries highlight the challenge of training on tasks that are intertwined, without clear delineation of when one task ends and another begins. Many existing methods fail to account for this characteristic, which is common in real-world learning scenarios. While there are other significant properties associated with continual learning in the real world, this study focuses on the often-neglected aspects of bounded memory and blurry task boundaries. By addressing these properties, we aim to develop methods that are more robust and applicable in practical settings.
@@ 2.2 Continual Learning Methods
Regularization methods in continual learning focus on incorporating constraints during model training to promote the retention of past knowledge. One simple approach is to include an $L_{2}$ penalty in the loss function. Elastic Weight Consolidation (EWC) builds upon this idea by assigning weights to parameters based on their importance for previous tasks using the Fisher information matrix (Kirkpatrick et al., 2017). MAS measures the sensitivity of parameter changes on the model’s output, prioritizing the retention of parameters with a larger effect (Aljundi et al., 2018). VCL leverages variational inference to minimize the Kullback-Leibler divergence between the current and prior parameter distributions (Nguyen et al., 2018). Progress and Compress learns new tasks using a separate model and subsequently distills this knowledge into the main model while safeguarding the previously acquired knowledge (Schwarz et al., 2018). However, regularization methods may struggle with blurry task boundaries as they rely on knowledge of task endpoints to apply regularization techniques effectively. In our experiments, EWC was chosen as the representative regularization method based on its performance in the original Continual World experiments (Wolczyk et al., 2021).
@@
Architecture-based methods aim to maintain distinct sets of parameters for each task, ensuring that future learning does not interfere with the knowledge acquired from previous tasks. Packnet (Mallya & Lazebnik, 2018), UCL (Ahn et al., 2019), and AGS-CL (Jung et al., 2020) all safeguard previous task information in a neural network by identifying important parameters and freeing up less important parameters for future learning. Identification of important parameters can be done through iterative pruning (Packnet), parameter uncertainty (UCL), and activation value (AGS-CL). However, a drawback of parameter isolation methods is that each task requires its own set of parameters, which may eventually exhaust the available parameters for new tasks and necessitate a dynamically expanding network without bounded memory (Yoon et al., 2018). Additionally, parameter isolation methods require training on a single task at a time to prune and isolate parameters, preventing concurrent learning from multiple interwoven tasks. In our experiments, PackNet was selected as the representative architecture-based method based on its performance in the original Continual World experiments (Wolczyk et al., 2021).
@@
Pseudo-rehearsal methods mitigate the forgetting of previous tasks by generating synthetic samples from past tasks and replaying them to the learner. Deep generative replay (DGR) (Figure 2) utilizes a generative model, such as generative adversarial networks (Goodfellow et al., 2020), variational autoencoders (Kingma & Welling, ), or diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015), to generate the synthetic samples. Originally, deep generative replay was proposed to address continual supervised learning problems, where the generator only needed to generate single data point samples (Shin et al., 2017). However, in decision-making tasks, expert demonstrations consist of trajectories (time-series) with a significantly higher-dimensional distribution space.
One existing DGR method generates individual state observations i.i.d. instead of entire trajectories. However, this approach leads to a higher sample complexity compared to generating entire trajectories. The sample complexity of generating enough individual state observations i.i.d. to cover every portion of the trajectory $m$ times can be described using the Double Dixie Cup problem (Newman, 1960). For trajectories of length $n$, it takes an average of $\Theta(n\log n+mn\log\log n)$ i.i.d. samples to ensure at least $m$ samples for each timestep. In scenarios with limited replay time (small $m$) and long trajectories (large $n$) the sample complexity can be approximated as $\Theta(n\log n)$ using the Coupon Collector’s problem (Myers & Wilf, 2006). The additional $\Theta(\log n)$ factor reduces the likelihood of achieving complete sample coverage of the trajectory when the number of replays or replay time is limited, especially considering the computationally expensive nature of current generative methods. Furthermore, there is a risk that the generator assigns different probabilities to each timestep of the trajectory, leading to a selective focus on certain timesteps rather than equal representation across the trajectory.
Another existing DGR method is autoregressive trajectory generation. In the existing autoregressive method, CRIL, a generator is used to generate samples of the initial state, and a dynamics model predicts the next state based on the current state and action (Gao et al., 2021). However, even with a dynamics model accuracy of 99% and a 1% probability of deviating from the desired trajectory, the probability of an autoregressively generated trajectory going off course is $1-0.99^{n}$, where $n$ denotes the trajectory length. With a trajectory length of $n=200$ (as used in our experiments), the probability of an autoregressively generated trajectory going off course is $0.871-0.99^{200}=$ 0.87. This example demonstrates how the issue of compounding error leads to a high probability of failure, even with a highly accurate dynamics model.
@@ 3.1 Imitation Learning
Imitation learning algorithms aim to learn a policy $\pi_{\theta}$ parameterized by $\theta$ by imitating a set of expert demonstrations $D=\{\tau_{i}\}_{i=1\ldots M}$. Each trajectory $\tau_{i}$ consists of a sequence of state-action pairs $\{(s_{j},a_{j})\}_{j=1\ldots|\tau_{i}|}$ where $|\tau_{i}|$ is the length of the trajectory. Each trajectory comes from a task $\mathcal{T}$ which is a Markov decision process that can be represented as a tuple $\langle S,A,T,\rho_{0}\rangle$ with state space $S$, action space $A$, transition dynamics $T:S\times A\times S\to[0,1]$, and initial state distribution $\rho_{0}$. Various algorithms exist for imitation learning, including behavioral cloning, GAIL (Ho & Ermon, 2016), and inverse reinforcement learning (Ng & Russell, 2000). In this work, we use behavioral cloning where the objective can be formulated as minimizing the loss function:
@@ 3.3 Diffusion Probabilistic Models
Diffusion probabilistic models (Ho et al., 2020; Sohl-Dickstein et al., 2015) generate data through a learned reverse denoising diffusion process $p_{\theta}(x_{t-1}\mid x_{t})$. The forward diffusion process $q(x_{t}\mid x_{t-1})$ gradually adds Gaussian noise to an input $x_{0}$ at each time step $t$, ultimately resulting in pure noise $x_{T}$ at $t=T$. The forward diffusion process is defined as:
@@
where $10<\beta_{t}<$ 1 is defined by a known variance schedule. In our implementation, we adopted the cosine schedule proposed by Nichol & Dhariwal (2021). For a sufficiently large time horizon $T$ and a well-behaved variance schedule, $x_{T}$ approximates an isotropic Gaussian distribution. If we had the reverse diffusion process $p(x_{t-1}\mid x_{t})$, we could sample $x_{T}\sim\mathcal{N}(0,\mathbf{I})$ and obtain a sample from $q(x_{0})$ by denoising $x_{T}$ with $p(x_{t-1}\mid x_{t})$. However, computing $p(x_{t-1}\mid x_{t})$ is intractable as it necessitates knowledge of the distribution of all possible $x_{t}$. Instead, we approximate $p(x_{t-1}\mid x_{t})$ using a neural network:
@@
Since $q$ and $p_{\theta}$ can be viewed as a variational auto-encoder (Kingma & Welling, ), we can use the variational lower bound to minimize the negative log-likelihood of the reverse process. We can express $\mu_{\theta}(x_{t},t)$ from Equation 4 as:
@@ 4 Method
We employ a U-net (Ronneberger et al., 2015) trained on the loss specified in Equation 6 to implement the generative diffusion model $G_{\gamma}$. Since we utilize proprioceptive observations in our experiments, $\pi_{\theta}$ is implemented with a multi-layer perceptron trained on the loss specified in Equation 1.
@@ 5.1 Experimental Setup
We evaluate our method on the Continual World benchmarks CW10 and CW20 (Wolczyk et al., 2021), along with our own variant of CW10 called BB10 that evaluates the ability of methods to handle blurry task boundaries. CW10 consists of a sequence of 10 Meta-World (Yu et al., 2020) tasks, where each task involves a Sawyer arm manipulating one or two objects in the Mujoco physics simulator. For computational efficiency, we provide the agents with proprioceptive observations. Notably, the observation and action spaces are continuous and remain consistent across all tasks. CW20 is an extension of CW10 with the tasks repeated twice. To our knowledge, Continual World is the only standard continual learning benchmark for decision-making tasks. BB10 gives data to the learner in 10 sequential buckets $B_{1},\ldots,B_{10}$. Data from task $\mathcal{T}_{i}$ from CW10 is split evenly between buckets $B_{i-1}$, $B_{i}$, and $B_{i+1}$, except for the first and last task. Task $\mathcal{T}_{1}$ is evenly split between buckets $B_{0}$ and $B_{1}$, and task $\mathcal{T}_{10}$ is evenly split between buckets $B_{9}$ and $B_{10}$.
In order to bound memory usage, task conditioning should utilize fixed-size embeddings like natural language task embeddings, rather than maintaining a separate final neural network layer for each individual task. For simplicity, we condition the model on the task with a one-hot vector as a proxy for natural language prompt embeddings. For BB10, the model is still conditioned on the underlying task rather than the bucket. Additionally, we do not allow separate biases for each task, as originally done in EWC (Kirkpatrick et al., 2017). Expert demonstrations for training are acquired by gathering 100 trajectories per task using hand-designed policies from Meta-World, with each trajectory limited to a maximum of 200 steps. Importantly, the learner model remains consistent across different methods and benchmark evaluations. Moreover, we maintain a consistent replay ratio of $r=0.9$ across all pseudo-rehearsal methods.